OpenStudy (anonymous):

a scientist prepares a solution by mixing 25ml of solution g with 50ml of solution h resulting in a solution with a mass of 125 grams . he then combined 45ml of solution g with 20 ml of solution h which procedures a solution with a mass of 120 grams in the equation below g represents the concentration of solution g and h represents the concentration of solution h . what is the concentration of solution h , in grams/ml

OpenStudy (anonymous):

Sure, basically if 25g +50h = 125 and 45g+ 20h = 120, You need to first solve one equation for either g or h. So 25g= 125 - 50h. g= 5-2h. Then you can plug in 5-2h into the other problem and get 45(5-2h) + 20h = 120.

OpenStudy (anonymous):

By solving for the variable in the first equation you can then substitut that information into the second.

OpenStudy (anonymous):

thank you